OpenStudy (anonymous):

Simplify 4/5 - 3/p will give metal to best answer A. 4p-15 over 5p B.1 over 5p C. - 9 over 5p D.- 9 over 5

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 } - \frac{ 3 }{ p } A. \frac{ 4p-15 }{ 5p } B. \frac{ 1 }{ 5p } C. -\frac{ 9 }{ 5p } D. -\frac{ 9 }{ 5 }\]

OpenStudy (lncognlto):

\[\frac{ 4 }{5 } - \frac{ 3 }{ p } = ?\]So, to get only one denominator, we multiply the 4/5 by p/p and the 3/p by 5/5.\[\left( \frac{ 4 }{ 5 } \times \frac{ p }{ p } \right) - \left( \frac{ 3 }{ p } \times \frac{ 5 }{5 } \right) = ?\]

OpenStudy (lncognlto):

Thus\[\frac{ 4p }{ 5p } - \frac{ 15 }{ 5p } = ?\] Thus the final answer is...?

OpenStudy (anonymous):

sorry, i have no clue

OpenStudy (anonymous):

A The answer is A

OpenStudy (lncognlto):

As the two fractions now have the same denominator, we can combine them into one: \[\frac{ 4p - 15 }{ 5p } = ?\]Thus \[\frac{ 4 }{ 5 } - \frac{ 3 }{ p } = \frac{ 4p - 15 }{ 5p }\]

OpenStudy (anonymous):

ok, thanks :)

OpenStudy (lncognlto):

@TheBadWolf I'm helping here, stop giving away the answer. Read the CoC of OS.